OpenStudy (anonymous):

Verify the identity. cos(4u) = cos^2(2u) - sin^2(2u)

OpenStudy (anonymous):

\[\cos \left( 4u\right) = \cos ^{2}\left( 2u \right) - \sin ^{2}\left( 2u \right)\]

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@arindameducationusc

OpenStudy (anonymous):

@kaitlyn_nicole

OpenStudy (anonymous):

@kaitlyn_nicole

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (arindameducationusc):

Do you know Cos2theta= cos^2theta-sin^2theta?

OpenStudy (anonymous):

yes, double angle identity

OpenStudy (michele_laino):

hint: we can rewrite the left side as below: \[\Large \cos \left( {4u} \right) = \cos \left( {2u + 2u} \right)\]

OpenStudy (michele_laino):

now you can apply the formula of addition for cosine function

OpenStudy (michele_laino):

\[\Large \cos \left( {x + y} \right) = \cos x\cos y - \sin x\sin y\]

OpenStudy (anonymous):

cos ( 2u + 2u ) = cos (2u) cos (2u) - sin (2u) sin (2u) ?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

what do i do from there?

OpenStudy (michele_laino):

it is simple, since: \[\Large \cos \left( {2u} \right)\cos \left( {2u} \right) = {\left( {\cos \left( {2u} \right)} \right)^2}\]

OpenStudy (michele_laino):

similarly for sin(2u)*sin(2u)

OpenStudy (anonymous):

sin(2u) sin(2u) = (sin(2u))^2 ?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

okay, so since im trying to verify this identity, what else do i do to get the answer to look like cos(4u) = cos(4u) ?

OpenStudy (michele_laino):

since we have proven that left side is equal to right side, then we are done here

OpenStudy (anonymous):

so is the final answer cos ( 2u + 2u ) = cos (2u) cos (2u) - sin (2u) sin (2u) or cos(2u) cos(2u) = (cos(2u))^2 ?

OpenStudy (michele_laino):

the answer is the entire procedure which shows how to get the right side, starting from left side

OpenStudy (anonymous):

ahhhhhh, i see, i see. thank you soooooo much for your help!

OpenStudy (michele_laino):

:)